OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = 2/3 and passing through (1, 1)

OpenStudy (anonymous):

Do you know the point-slope formula?

OpenStudy (anonymous):

yes. i got 3x-2?

OpenStudy (anonymous):

its wrong. correct one is '2x-3y+1=0'

OpenStudy (anonymous):

No its not chenna. In standard form it would be 2x-3y=-1.

OpenStudy (anonymous):

syandard form is y=(2/3)x+1/3

OpenStudy (anonymous):

remember y=mx+c

OpenStudy (anonymous):

mlhuyck, what you would do is take: \[y-y_0=m(x-x_0)\] Then (1,1) would be your (x0,y0) So you have: \[y-1=\frac{2}{3}(x-1)\] \[y-1=\frac{2}{3}x-\frac{2}{3}\] \[y-\frac{1}{3}=\frac{2}{3}x\] \[-\frac{1}{3}=\frac{2}{3}x-y\] Clean it up by multiplying by 3. \[-1=2x-3y\]

OpenStudy (anonymous):

Standard form is \[Ax+By=C\]

OpenStudy (anonymous):

Does that make more sense mlhuyck?

OpenStudy (anonymous):

but the standard form for a straight line equation is y=mx+c and its a fact

OpenStudy (anonymous):

No it is not. That is slope intercept form.

OpenStudy (anonymous):

Definitions: Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

OpenStudy (anonymous):

y=mx+c there m is the gradient and c is the intercept

OpenStudy (anonymous):

Hashir. That is slope intercept form. And its typically y=mx+b where b is the y intercept and m is the slope. I will bet you any amount of money.

OpenStudy (anonymous):

i m sure of it i have solved more than 1000 question with thaat fact

OpenStudy (anonymous):

but books are more authentic than this sites

OpenStudy (anonymous):

Do you want me to scan my algebra 1 textbook?

OpenStudy (anonymous):

One of those websites is the personal website of an instructor at Lousiana University.